# Dysca: A Dynamic and Scalable Benchmark for Evaluating Perception Ability of LVLMs

### v2

@@ 3.2 Perceptual Tasks
Evaluation dimensions. Perception is one of the most fundamental capabilities of LVLMs and previous works [25] have shown that the lack of perceptual ability may result in hallucination [57]. In order to comprehensively evaluate LVLMs’ perception capability, we design 20 perceptual subtasks where we show all subtasks and the corresponding amount of their annotation in the Fig. 2. We investigate on two types of perception dimensions, i.e., coarse-grained and fine-grained perception. Coarse-grained perception involves recognizing the style, background and color of images. Fine-grained perception involves recognizing the animal, object, plant, food, age, gender, expression, race, celebrity, action, text, clothes, movie, anime, landmark, profession and TV shows.
Data sources. For each perceptual subtask, we collect the textual data first to construct the metadata $M$. For the TV shows, Anime and Movie, we select the titles from the rating list of IMDb***https://www.imdb.com/ based on the number of user reviews. For the styles, we utilize the style lists collected from the communityhttps://stable-diffusion-art.com/sdxl-styles/ and remove those which have strong reflect on the image content like "architectural style" and "Pokemon style". Note that the style list does not include the style prompt associated with a particular artist’s name. Besides, for the remaining contents, we select them from the label of current dataset (e.g., ImageNet [54]). All the selected textual data above constitute the metadata $M$. We provide the detailed information of the metadata in the Appendix C.
@@ 3.3 Construction of Questions & Answers
Recall that the data generation for Dysca benchmark consists of four dimensions, i.e., $(M,P,I,Q)$, denoting the metadata ($M$), prompt ($P$), image ($I$) and question-answer pairs ($Q$), respectively. The relationships between these parts and the process of constructing Dysca are shown in Fig. 3. The metadata M is the core of the whole Dysca, containing all the information for generating $P$, $I$ and $Q$. The metadata $M$ consists of foreground, attribute, background and style, and these information guide the generation of the prompt ($P$) through pre-designed templates. Then, we utilize the T2I diffusion model to generate the corresponding image using the prompt $P$. For generating the image with specific text on it for the OCR subtask, we leverage TextDiffusion2 [58], which is the state-of-the-art text rendering method. For the rest of images, we leverage Stable Diffusion XL [36]. Subsequently, based on the different question types we select, i.e., multi-choices, true-or-false and free-form, we generate the corresponding Vision-language QA pairs in Dysca.
@@
Besides, in order to evaluate the model performance under various scenarios, we conduct experiments on 4 scenarios, i.e., clean, corruption, print attacking and adversarial attacking. For the print attacking, followed by [35], we add the deceptive text on the image, where the text is a wrong option. Besides, to comprehensively evaluate the performance of LVLMs under corruption scenario, we add more typographic factors to original settings (i.e., different font orientations and font positions). For the adversarial attacking, we leverage PGD [59] to generate the adversarial image. We use InstructBLIP [3] as the proxy model and regard others as the black box models. The reason why we choose InstructBLIP is that it has shown superior performance in clean scenario. Besides, the black-box setting better reflects the robustness of the models when they face the real-world adversarial attacks. For the corruption, we leverage the image corruption methods collected from [34]. We remove some hard corruptions as they significantly impact the quality of the image, leading to human failure in judging the style and content of the image. The detailed examples are shown in Appendix D.
@@
Consider that the Text-to-image diffusion model may generate the failure cases that affect the quality of the proposed benchmark, we leverage the off-the-shelf models, i.e., PP-OCRv3 [56] and CLIP-L-14 [16], to clean the data. PP-OCRv3 [56] is leveraged as the filter to exclude the failure image that TextDiffusion2 [58] generates the wrong text on the image. For the other images, we use CLIP-L-14 [16] to filter out the images with low text-image consistency. In the end, we filter out nearly 15% of low quality samples. The final statistics of our released Dysca are shown in Tab. 2. Note that the OCR subtask does not involve print attacking scenario as misidentifying adversarial text does not indicate poor OCR robustness of the LVLMs. Therefore, there are 7K fewer questions in the print attacking scenario. Besides, for the free-form question type, since it allows to assess the model’s perception abilities across multiple subtasks at the same time, we reduce the number of free-form questions for achieving a balanced data distribution.
@@ 3.4 Evaluation Strategy
Table 3: Evaluation results on the 20 perceptual subtasks. The top two results on each subtask are bolded and underlined, respectively. "MC" and "TF" indicate the accuracy (%) of "Multi-choices" and "True-or-false", respectively.
Movie Action TV Show Profession Landmark
Model Language Model MC TF MC TF MC TF MC TF MC TF
Blip2 [1] Flan-T5-XL 71.46 68.61 96.24 93.04 56.29 61.99 77.88 75.87 98.70 94.58
InstructBLIP [3] Flan-T5-XL 77.35 67.38 97.28 93.61 65.89 60.96 79.17 76.51 98.04 95.75
XComposer-VL [7] InternLM-7B 81.90 78.36 97.03 94.75 77.81 73.29 83.97 75.87 98.04 94.81
LLava-1.5 [48] Vicuna-7B 55.95 51.68 74.78 50.75 56.29 50.68 58.65 56.83 85.00 54.48
LLava-1.5 [48] Vicuna-13B 66.29 54.04 77.93 59.71 59.27 54.11 66.67 59.05 94.35 55.90
MiniGPT-4 [2] Vicuna-7B 32.68 52.35 42.96 51.00 30.46 50.00 33.97 50.16 53.26 52.36
Otter [5] LLaMA-7B 65.36 59.08 65.85 72.36 68.54 57.53 66.99 54.92 58.48 66.51
Qwen-VL-Chat [49] Qwen-7B 71.25 51.01 96.35 63.14 67.22 49.32 75.64 54.60 96.96 54.72
Shikra [6] Vicuna-7B 66.29 59.64 78.00 77.29 60.26 56.85 78.85 68.25 89.35 70.99
Shikra-VQA [6] Vicuna-7B 66.39 61.66 96.17 80.32 59.93 60.27 78.21 67.94 92.83 72.88
Anime Clothes Celebrity Food Plant
Model Language Model MC TF MC TF MC TF MC TF MC TF
Blip2 [1] Flan-T5-XL 57.07 61.94 82.38 73.79 81.98 76.48 91.52 88.04 92.50 91.19
InstructBLIP [3] Flan-T5-XL 61.28 64.04 86.51 81.08 83.69 76.14 91.71 88.85 93.31 91.92
XComposer-VL [7] InternLM-7B 75.41 74.58 86.18 85.73 88.53 87.28 92.33 90.21 93.25 93.14
LLava-1.5 [48] Vicuna-7B 47.15 48.46 47.19 50.41 57.36 54.72 54.63 51.10 49.08 48.59
LLava-1.5 [48] Vicuna-13B 57.61 50.14 65.20 57.20 60.14 57.94 78.87 57.45 79.72 55.51
MiniGPT-4 [2] Vicuna-7B 29.21 48.74 31.25 50.47 28.62 49.64 44.66 50.78 45.38 50.53
Otter [5] LLaMA-7B 61.82 57.30 47.13 69.52 42.41 63.52 46.81 79.62 66.41 81.51
Qwen-VL-Chat [49] Qwen-7B 71.60 54.63 78.44 60.72 87.37 50.05 89.37 53.37 92.32 59.91
Shikra [6] Vicuna-7B 47.96 57.72 75.21 59.65 63.76 59.86 83.30 68.37 88.55 66.60
Shikra-VQA [6] Vicuna-7B 49.05 57.58 76.57 63.10 63.73 62.13 89.25 71.22 88.34 71.76
Age Gender Expression Race Animal
Model Language Model MC TF MC TF MC TF MC TF MC TF
Blip2 [1] Flan-T5-XL 62.61 59.65 99.37 94.86 89.27 75.92 74.38 71.95 96.64 95.27
InstructBLIP [3] Flan-T5-XL 65.14 59.55 99.51 88.21 91.85 79.34 77.99 74.62 97.21 94.98
XComposer-VL [7] InternLM-7B 67.89 78.35 99.60 98.06 89.64 82.52 80.42 76.69 97.83 96.70
LLava-1.5 [48] Vicuna-7B 38.35 55.15 54.14 49.78 63.86 49.76 43.98 50.56 49.26 50.90
LLava-1.5 [48] Vicuna-13B 49.55 59.59 98.71 83.53 71.29 58.33 70.16 62.84 85.99 58.23
MiniGPT-4 [2] Vicuna-7B 31.75 51.71 56.49 49.22 42.01 50.91 27.54 50.41 45.97 49.57
Otter [5] LLaMA-7B 37.98 51.19 78.23 77.96 73.86 60.36 43.00 57.88 81.32 83.63
Qwen-VL-Chat [49] Qwen-7B 53.48 49.09 97.72 58.55 85.22 64.30 73.50 56.88 95.60 64.37
Shikra [6] Vicuna-7B 65.37 56.71 97.85 73.16 90.47 70.55 74.17 54.88 89.82 69.06
Shikra-VQA [6] Vicuna-7B 65.50 57.34 99.39 81.41 92.36 74.71 73.90 55.24 91.18 74.26
Object OCR Style Background Color
Model Language Model MC TF MC TF MC TF MC TF MC TF
Blip2 [1] Flan-T5-XL 89.32 87.75 71.89 62.07 99.86 88.62 63.97 68.37 88.01 85.55
InstructBLIP [3] Flan-T5-XL 89.87 90.13 72.61 60.93 99.96 81.23 66.50 69.19 90.93 85.55
XComposer-VL [7] InternLM-7B 90.23 91.72 69.22 78.66 95.83 81.32 71.26 75.27 87.97 89.02
LLava-1.5 [48] Vicuna-7B 58.63 49.99 47.08 51.02 46.35 50.49 44.33 50.95 41.91 51.90
LLava-1.5 [48] Vicuna-13B 77.86 57.93 58.19 53.17 85.80 51.56 62.67 55.33 63.81 52.90
MiniGPT-4 [2] Vicuna-7B 51.75 51.55 31.37 51.56 43.41 48.64 31.94 50.10 35.29 50.51
Otter [5] LLaMA-7B 47.31 82.51 66.23 61.41 98.04 61.58 63.06 63.40 48.28 57.97
Qwen-VL-Chat [49] Qwen-7B 88.32 63.14 71.74 51.86 98.21 52.43 69.42 51.59 86.39 53.05
Shikra [6] Vicuna-7B 70.91 69.22 60.79 53.63 95.13 58.95 71.24 62.66 84.91 61.83
Shikra-VQA [6] Vicuna-7B 89.43 76.00 59.63 54.14 99.51 61.48 70.98 65.96 83.79 64.49
@@ 4 Results and Analysis
In this section, we report the evaluation results and make insightful analysis. A total of 8 LVLMs with 10 checkpoints are evaluated on Dysca benchmark, including BLIP2 [1], InstructBLIP [3], LLavA [48], MiniGPT-4 [2], Otter [5], XComposer-VL [7], Qwen-VL-chat [49] and Shikra [6]. Each model is evaluated with all the 20 perception subtasks under 4 scenarios. The detailed rankings for each subtask can be found in the Appendix A.
@@ 4.1 Main Results
Clean Scenario. The evaluation results of various LVLMs in different perceptual subtasks under clean scenarios are presented in Tab. 3. Since the evaluation for free-form question type usually involves multiple subtasks, we can not calculate the results of free-form for each subtask individually. Instead, we display the overall score of free-form in the first row in Tab. 4. As can be seen, Xcomposer-VL [7] outperforms other LVLMs, achieving top-1 or top-2 results in most subtasks, but InstructBLIP [3], Qwen-VL-chat [49] and BLIP2 also take a lead in a few subtasks.
Noisy Scenarios The evaluation results of various LVLMs under noisy scenarios (i.e., corruption, print attacking and adversarial attacking) are presented in Tab. 4. As can be seen, For the multi-choices and true-or-false question type, Xcomposer-VL [7] still takes a lead on all 4 scenarios. For the free-form, LLava-1.5-7b [48] achieves the best.
Table 4: Evaluation results on 4 scenarios. "MC", "TF" and "FF" indicates "Multi-choices", "True-or-false" and "Free-form", respectively. "PrintAtt" and "AdverAtt" means "Print Attacking" and "Adversarial Attacking", respectively. "*": the model is under white-box setting.
Scenarios Blip2 [1] InstructBLIP [3] XComposer-VL [7] LLava-1.5-13b [48] LLava-1.5-7b [48]
MC TF FF MC TF FF MC TF FF MC TF FF MC TF FF
Clean 82.07 78.78 37.56 84.29 79.00 38.94 86.22 84.82 46.03 71.50 57.72 48.61 53.70 51.41 50.15
Corruption 81.17 77.98 37.32 83.77 78.57 38.72 85.20 84.16 45.52 70.34 57.57 48.28 53.21 51.40 49.54
PrintAtt 64.78 68.01 35.82 59.18 58.80 35.71 74.08 72.71 44.46 48.91 54.72 48.47 40.02 50.94 48.38
AdverAtt 23.28 48.98 25.70 21.49* 48.74* 29.15* 20.85 49.54 21.17 66.86 56.56 47.11 50.70 51.36 48.17
Scenarios MiniGPT-4 [2] Otter [5] Qwen-VL-Chat [49] Shikra [6] Shikra-VQA [6]
MC TF FF MC TF FF MC TF FF MC TF FF MC TF FF
Clean 38.50 50.51 38.90 61.36 65.99 39.12 82.31 55.84 49.90 76.61 63.79 47.17 79.31 66.69 29.93
Corruption 38.70 49.99 40.35 62.19 65.36 38.71 79.26 52.74 49.74 76.42 64.31 46.43 79.06 66.95 29.40
PrintAtt 36.16 50.00 42.45 38.11 36.27 36.93 59.68 46.40 48.65 60.02 40.18 47.03 62.20 41.05 27.91
AdverAtt 26.75 49.52 22.59 57.78 62.27 35.76 78.47 58.04 44.67 72.20 62.16 43.56 74.04 64.64 26.81
4.2 Analysis
4.2.1 Key Observations
(1) For individual models, their perceptual performance varies across different subtasks. For example, in the case of Qwen-VL-Chat [49], it achieves a score of 96.96% accuracy in landmark recognition task for multi-choices questions (2% below the first-place score), but obtains a score of 53.48% accuracy in age recognition task (12% below the first-place score). The results suggest that Qwen-VL-Chat [49] may require more fine-tuning in age perception data. Analyzing the performance of the models across various subtasks contribute to purposive improving.
Figure 4: Models exhibit different performance when facing the same image but different question types.
(2) Models exhibit performance inconsistency when facing multi-choices and true-or-false question types. As can be seen, in the object recognition subtask, Otter [5] achieves an accuracy of 47.31% in the multi-choices question type (22% higher than random guessing), but obtains an accuracy of 82.51% in the true-or-false question type (32% higher than random guessing). Interestingly, we observe the opposite results in the Qwen-VL-Chat [49]. In the object recognition subtask, it achieves an accuracy of 88.32% in the multi-choices but achieves an accuracy of 63.14% in the true-of-false. We also observe the same problem shown in other models and we display two examples in Fig. 4. We speculate that the inconsistency may be attributed to the bias in the training dataset towards particular question types, such as using more multi-choices or true-or-false questions.
(3) Each model exhibits robustness in the corruption scenario, but suffers degradation in the two attacking scenarios. As shown, all models exhibit minor score variation of less than 1% under the corruption scenario. However, they exhibit degradation when facing print attacking (e.g., 84.29% vs. 59.18% for InstructBlip [3] in multi-choices accuracy). XComposer-VL [7] shows the strongest robustness, maintaining over 70% accuracy for both multi-choices and true-or-false. Besides, since our adversarial algorithm specifically targets the image encoder, the LVLMs that share the same encoder architecture (i.e., Blip2, InstructBLip and XComposer-VL all using EVA-CLIP [62] as the image encoder) exhibit significant performance degradation, with accuracy even falling below random selection. Models utilizing alternative image encoders also experience a performance decrease of approximately 5% to 10%. A more detailed result can be found in Appendix D.
4.2.2 The Validity of Dysca

### v3

@@ 3.2 Perceptual Tasks
Evaluation dimensions. Perception is one of the most fundamental capabilities of LVLMs and previous works (Fu et al., 2023) have shown that the lack of perceptual ability may result in hallucination (Li et al., 2023e). In order to comprehensively evaluate LVLMs’ perception capability, we collect and organize existing sub-dimensions from current evaluation datasets, resulting in a total of 20 assessment dimensions where we show all the subtasks and the corresponding amount of their annotation in the Fig. 2. We investigate on two types of perception dimensions, i.e., coarse-grained and fine-grained perception. Coarse-grained perception involves recognizing the style, background and color of images. Fine-grained perception involves recognizing the animal, object, plant, food, age, gender, expression, race, celebrity, action, text, clothes, movie, anime, landmark, profession and TV shows.
Data sources. For each perceptual subtask, we collect the textual data first to construct the metadata $M$. For the TV shows, Anime and Movie, we select the titles from the rating list of IMDb [^1] https://www.imdb.com/ based on the number of user reviews. For the styles, we utilize the style lists collected from the community [^2] https://stable-diffusion-art.com/sdxl-styles/ and remove those which have strong reflect on the image content like “architectural style” and “Pokemon style”. Note that the style list does not include the style prompt associated with a particular artist’s name. Besides, for the remaining contents, we select them from the label of current dataset (e.g., ImageNet (Russakovsky et al., 2014)). All the selected textual data above constitute the metadata $M$. We provide the detailed information of the metadata in the Appendix E.
@@ 3.3 Construction of Questions & Answers
Recall that the data generation for Dysca benchmark consists of four dimensions, i.e., $(M,P,I,Q)$, denoting the metadata ($M$), prompt ($P$), image ($I$) and question-answer pairs ($Q$), respectively. The relationships between these parts and the process of constructing Dysca are shown in Fig. 3. The metadata M is the core of the whole Dysca, containing all the information for generating $P$, $I$ and $Q$. The metadata $M$ consists of foreground, attribute, background and style, and these information guide the generation of the prompt ($P$) through pre-designed templates. Then, we utilize the T2I diffusion model to generate the corresponding image using the prompt $P$. For generating the image with specific text on it for the OCR subtask, we leverage TextDiffusion2 (Chen et al., 2023a), which is the state-of-the-art text rendering method. For the rest of images, we leverage Stable Diffusion XL (Podell et al., 2023). Subsequently, based on the different question types we select, i.e., multi-choices, true-or-false and free-form, we generate the corresponding VQA pairs in Dysca.
@@
Besides, in order to evaluate the model performance under various scenarios, we conduct experiments on 4 scenarios, i.e., clean, corruption, print attacking and adversarial attacking. For the print attacking, followed by (Cheng et al., 2024), we add the deceptive text on the image, where the text is a wrong option. Besides, to comprehensively evaluate the performance of LVLMs under corruption scenario, we add more typographic factors to original settings (i.e., different font orientations and font positions). For the adversarial attacking, we leverage PGD (Madry et al., 2017) to generate the adversarial image. We use InstructBLIP (Dai et al., 2023) as the proxy model and regard others as the black box models. The reason why we choose InstructBLIP is that it has shown superior performance in clean scenario. Besides, the black-box setting better reflects the robustness of the models when they face the real-world adversarial attacks. For the corruption, we leverage the image corruption methods collected from (Zhang et al., 2024b). We remove some hard corruptions as they significantly impact the quality of the image, leading to human failure in judging the style and content of the image. The detailed examples are shown in Appendix F.
@@
Data Clean. To ensure the quality of Dysca, four steps are adopt: 1) First, we manually remove difficult-to-generate foregrounds and attributes, along with backgrounds and styles that could heavily affect image content. We believe this process can serve as a coarse-grained method to eliminate samples that are highly likely to be generated incorrectly. 2) Then, we leverage the off-the-shelf models, i.e., PP-OCRv3 (Li et al., 2022) and CLIP-L-14 (Radford et al., 2021), to clean the data. PP-OCRv3 (Li et al., 2022) is leveraged as the filter to exclude the failure image that TextDiffusion2 (Chen et al., 2023a) generates the wrong text on the image. For the other images, we use CLIP-L-14 (Radford et al., 2021) with a threshold of 0.75 to filter out the images with low text-image consistency. We find that using 0.75 as the threshold achieves a good balance between image correctness and data scale. 3) After that, We select the top six performing models and eliminate any question-answer pairs where the models either answer incorrectly or indicate that the answer was not included among the options. We observe that nearly 100% of the samples filtered out by these models are incorrect. 4) Finally, we analyze the patterns in these incorrect samples, removing the associated vocabulary from our metadata and discarding all related samples. By meticulously refining the metadata manually and utilizing automated tools to assist in question filtering, Dysca ensures high-quality data synthesis. In the end, we filter out nearly 40% of low quality samples. The final statistics of our released Dysca are shown in Tab. 2. Note that the OCR subtask does not involve print attacking scenario as misidentifying adversarial text does not indicate poor OCR robustness of the LVLMs. Therefore, there are 7K fewer questions in the print attacking scenario. Besides, for the free-form question type, since it allows to assess the model’s perception abilities across multiple subtasks at the same time, we reduce the number of free-form questions for achieving a balanced data distribution.
@@ 3.4 Evaluation Strategy
Figure 4: The failure cases for the noisy scenarios. From left to right are: corruption scenario, adversarial attacking scenario, and print attacking scenario.
@@
Besides, thanks to generative evaluation framework of Dysca, we are able to effectively control variables and conduct a detailed analysis of the model’s fine-grained capabilities. Specifically, we introduce two novel metrics to measure the sensitivity of LVLMs on question types and covariate shift. The sensitivity to question type aims to evaluate whether LVLMs exhibit inconsistent performance when facing different types of question types (i.e., multi-choice vs. true-or-false). We first normalize the score of true-or-false by $TFSQ=\frac{S-50}{100-50}*100\%$ and multi-choices by $MCSQ=\frac{S-25}{100-25}*100\%$, where $S$ denotes the score of the models in each question type. The sensitivity of LVLMs to question types is defined by:
$SQ=\frac{(TFSQ-SQ_{Avg})^{2}+(MCSQ-SQ_{Avg})^{2}}{2},$ (1)
where $SQ_{Avg}=\frac{TFSQ+MCSQ}{2}$.
The sensitivity to covariate shift aims to evaluate whether LVLMs exhibit inconsistent performance when facing the same content and question format, but with variations in image covariates (i.e., image style). It is defined by:
$SC=\frac{\sum_{i=1}^{N}(S_{i}-SC_{Avg})^{2}}{N},$ (2)
where $SC_{Avg}=\frac{\sum_{i=1}^{N}S}{N}$ and $S_{i}$ denotes the score of the models in each style. Since there are 51 image styles in Dysca, we set $N=51$.
@@ 4 Results and Analysis
In this section, we report the evaluation results and make insightful analysis. A total of 26 LVLMs are evaluated on Dysca, including BLIP2 (Li et al., 2023d), InstructBLIP (Dai et al., 2023), LLavA (Liu et al., 2023a), MiniGPT-4 (Zhu et al., 2023), Otter (Li et al., 2023a), XComposer-VL (Zhang et al., 2023), Qwen-VL-chat (Bai et al., 2023a), Shikra (Chen et al., 2023b), Emu2-Chat (Sun et al., 2024), GLM-4V (GLM et al., 2024), MiniCPM-v2.5 (Yao et al., 2024), Yi-VL (AI et al., 2024), mPLUG-Owl-2 (Ye et al., 2023), Phi-3-Vision (Abdin et al., 2024), GPT-4o (OpenAI, 2023), Gemini-1.5-pro (Team et al., 2024). Each model is evaluated with all the 20 perception subtasks under 4 scenarios. The detailed rankings for each subtask can be found in the Appendix A.
@@ 4.1 Main Results
Blind Setting. We first evaluate LVLMs when only textual questions are provided. As shown in the “Blind” column of Tab. 3, all LVLMs yield consistent results on the Dysca and perform comparable to random guessing. This outcome demonstrates that the generated paradigm employed by Dysca effectively mitigates the potential impact of data leakage  (Chen et al., 2024), thereby enhancing the fairness of the evaluation results. Additional comparisons can be found in Appendix D.
Clean Scenario. The evaluation results of various LVLMs in different perceptual subtasks under clean scenarios are presented in the “clean” column of Tab. 3. We calculate the average score of 3 question types. As can be seen, GLM-4V (GLM et al., 2024) outperforms other LVLMs, achieving top-1 performance. MiniCPM-v2.5 (Yao et al., 2024), Xcomposer2-VL (Dong et al., 2024) and Gemini-1.5-pro (Team et al., 2024) also perform well. It is evident that for the latest large models, their scores remain below 85. The results highlight that all existing LVLMs still struggle to provide accurate responses to questions formulated by Dysca.
Noisy Scenarios. The evaluation results of various LVLMs under noisy scenarios (i.e., corruption, print attacking and adversarial attacking) are presented in last 3 columns in Tab. 3. The value in the brackets shows the relative values with respect to the ones in the clean scenario. As can be seen, GLM-4V (GLM et al., 2024) still takes a lead on corruption and adversarial attacking scenarios. For the print attacking scenario, mPLUG-Owl-2 (Ye et al., 2023) performs the best. Here, we present a failure case sample for each of the three different scenarios on Fig. 4.
Table 3: Evaluation results of 26 LVLMs, where the darker colors represent better performance. The top 1 result on each column are bolded and the value in brackets is the relative values with respect to the ones in the clean scenario. “PrintAtt” and “AdverAtt” means “Print Attacking” and “Adversarial Attacking”, respectively. “*”: the model is under white-box setting.
Scenarios
Model LLM Visual Encoder Blind Clean $\uparrow$ Corruption $\uparrow$ AdverAtt $\uparrow$ PrintAtt $\uparrow$ SQ $\downarrow$ SC $\downarrow$
MiniGPT-4 Vicuna-7B EVA-CLIP ViT-G 35.37 41.38 42.30 (+0.92) 34.42 (-6.96) 42.71 (+1.33) 193.6 1.7
MiniGPT-4 Vicuna-13B EVA-CLIP ViT-G 35.21 50.17 49.63 (-0.54) 31.77 (-18.40) 47.55 (-2.62) 758.0 2.5
MiniGPT-4 LLaMA2-7B EVA-CLIP ViT-G 34.77 56.61 55.70 (-0.91) 33.55 (-23.06) 49.78 (-6.83) 1344.5 3.0
MiniGPT-2 LLaMA2-7B EVA-CLIP ViT-G 35.28 58.46 58.06 (-0.40) 56.62 (-1.84) 52.96 (-5.50) 1512.1 4.3
BLIP2 Flan-T5-XL EVA-CLIP ViT-G 35.35 65.30 66.09 (+0.79) 32.55 (-32.75) 57.01 (-8.29) 165.5 8.5
BLIP2 OPT-3B EVA-CLIP ViT-G 34.99 39.54 40.29 (+0.75) 30.62 (-8.92) 37.26 (-2.28) 110.7 2.3
BLIP2 OPT-7B EVA-CLIP ViT-G 35.21 39.55 41.12 (+1.57) 31.76 (-7.79) 38.82 (-0.73) 50.2 1.8
InstructBLIP Vicuna-7B EVA-CLIP ViT-G 35.14 67.54 67.01 (-0.53) 34.42 (-33.12) 52.58 (-14.96) 1287.1 2.8
InstructBLIP Vicuna-13B EVA-CLIP ViT-G 34.37 64.89 64.68 (-0.21) 31.77 (-33.12) 53.53 (-11.36) 1252.4 2.5
InstructBLIP Flan-T5-XL EVA-CLIP ViT-G 34.51 66.54 67.58 (+1.04) 32.95 (-33.59)* 52.09 (-14.45) 271.3 6.3
InstructBLIP Flan-T5-XXL EVA-CLIP ViT-G 34.82 68.65 69.79 (+1.14) 32.95 (-35.70) 57.73 (-10.92) 215.0 6.8
LLava-1.5 Vicuna-7B CLIP ViT-L 34.63 51.27 51.70 (+0.43) 49.62 (-1.65) 47.27 (-4.00) 788.1 3.2
LLava-1.5 Vicuna-13B CLIP ViT-L 35.21 59.23 59.58 (+0.35) 56.87 (-2.36) 51.69 (-7.54) 912.5 7.7
Otter LLaMA-7B CLIP ViT-L 35.19 54.90 56.02 (+1.12) 51.42 (-3.48) 37.78 (-17.12) 427.4 6.0
Shikra LLaMA-7B CLIP ViT-L 34.96 62.24 63.06 (+0.82) 58.78 (-3.46) 49.56 (-12.68) 440.6 5.0
Xcomposer-VL InternLM-7B EVA-CLIP ViT-G 32.33 71.40 72.08 (+0.68) 30.28 (-41.12) 64.71 (-6.69) 147.9 6.1
Xcomposer2-VL InternLM2-7B CLIP ViT-L 32.76 79.13 78.64 (-0.49) 76.60 (-2.53) 66.34 (-12.79) 20.9 1.9
Qwen-VL-Chat Qwen-7B OpenClip ViT-bigG 33.06 62.18 61.05 (-1.13) 59.85 (-2.33) 51.94 (-10.24) 1885.4 6.7
Emu2-Chat LLaMA-33B EVA2-CLIP-E 35.14 63.64 62.81 (-0.83) 61.90 (-1.74) 54.82 (-8.82) 2497.9 1.6
GLM-4V GLM-4-9B-Chat EVA2-CLIP-E 35.08 82.09 81.95 (-0.14) 80.72 (-1.37) 52.09 (-30.00) 25.5 0.8
MiniCPM-V2.5 Llama3-Instruct 8B SigLIP SoViT-400m 34.99 78.75 77.41 (-1.34) 75.44 (-3.31) 60.77 (-17.98) 38.3 1.7
Yi-VL Yi-6B-Chat OpenClip ViT-H 35.01 75.71 74.94 (-0.77) 72.53 (-3.18) 64.97 (-10.74) 233.1 1.9
mPLUG-Owl-2 LLaMA2-7B CLIP ViT-L 35.03 74.09 72.85 (-1.24) 69.76 (-4.33) 72.85 (-1.24) 180.9 2.2
Phi-3-Vision Phi-3 CLIP ViT-L 34.74 73.23 72.11 (-1.12) 69.66 (-3.57) 57.78 (-15.45) 292.7 1.5
GPT-4o / / 35.02 75.69 75.52 (-0.17) 73.47 (-2.22) 56.34 (-19.35) 67.4 1.9
Gemini-1.5-Pro / / 34.55 77.79 77.12 (-0.67) 75.89 (-1.90) 61.05 (-16.74) 439.6 1.6
4.2 Key Observations
(1) For LVLMs, the capacity of the language model plays a crucial role. When using the same visual encoder, models that adopt a language model with larger parameter sizes (e.g., Vicuna-13B models generally outperform Vicuna-7B models by 8%) or models with a different but stronger architecture (e.g., GLM-4V (GLM et al., 2024) with GLM-4-9B-Chat shows a 20% performance increase compared to Emu2 (Sun et al., 2024) with LLaMA-33B) tend to achieve better performance.
(2) Each model shows robustness in the corruption scenario, but experiences significant degradation in both attack scenarios. In the image corruption scenario, all models demonstrate minimal score variations (less than 1%), indicating consistent performance under non-targeted disruptions. However, under print attacks, performance drops are notable. For example, two closed-source models experience substantial declines: Gemini-1.5-pro (Team et al., 2024) drops by 21.5% (from 77.79 to 61.05), and GPT-4o (OpenAI, 2023) suffers a 25.8% decrease (from 75.69 to 56.10). Among the advanced open-source models, GLM-4V’s (GLM et al., 2024) performance decreases sharply by 36.5% (from 82.09 to 52.09), and Phi-3-Vision (Abdin et al., 2024) records a 21.39% drop (from 73.23 to 57.78). Notably, mPLUG-Owl-2 (Ye et al., 2023) shows the highest robustness, with only a 1.7% reduction (from 74.09 to 72.84), while the XComposer-VL (Zhang et al., 2023; Dong et al., 2024) series also exhibits strong resistance to print attacks. In the adversarial attack scenario, where the attack algorithm directly targets the image encoder, LVLMs with shared encoder architecture (e.g., Blip2 (Li et al., 2023d), InstructBLIP (Dai et al., 2023), and XComposer-VL (Zhang et al., 2023), all using EVA-CLIP (Fang et al., 2022) as their image encoder) suffer significant performance drops, with some models even performing worse than random chance. For instance, XComposer-VL (Zhang et al., 2023) experiences a 57.6% drop (from 71.40 to 30.28). Models with alternative image encoders also experience degradation, ranging from 1% to 5%. The greater impact of adversarial noise compared to corruption noise suggests that adversarial attacks exhibit a certain degree of transferability across different model architectures. More detailed results are available in Appendix F.
Figure 5: The Visualization of hierarchical clustering results of 20 subtasks.
Figure 6: The Visualization of hierarchical clustering results of 51 animal categories.
(3) Models exhibit varying sensitivity to different question types and covariate shifts. For the Sensitivity to Question Types, Xcomposer2-VL (Dong et al., 2024) achieves the best result with a score of 20.9. However, we observe that the perception ability of LVLM does not show a positive correlation with sensitivity to question types. For instance, while the GLM-4V (GLM et al., 2024) achieves the highest performance in evaluation tasks, it exhibits higher sensitivity to question types than Xcomposer2-VL (Dong et al., 2024). One of the factors influencing the sensitivity to question types may be the inherent biases of the language model. Using the same base language model may result in similar outcomes for this metric. it is also noted that Gemini (Team et al., 2024) performs bad in this metric, revealing its preference for certain question types. For the Sensitivity to Covariate Shifts, GLM-4V (Dong et al., 2024) achieves the best result with a score of 0.8. However, we also observe that the perception ability of LVLM does not show a positive correlation with covariate shifts. For example, InstructBLIP-Flan-T5-XXL (Dai et al., 2023) outperforms InstructBLIP-Flan-T5-XL in terms of performance but shows higher sensitivity to covariate shifts.
4.3 Analysis on inter-task and intra-task
Analysis on Inter-task. Here, we perform hierarchical clustering based on the Euclidean distance of the scores across 20 evaluation dimensions for 26 models. As shown in the Fig. 6, based on the consistency of model performance across the 20 evaluation dimensions, the models can be categorized into four tiers. We also observe that LVLMs tend to perform better in dimensions that involve well-defined image perception, such as landmark recognition and object recognition. However, dimensions likes style recognition and movie recognition exhibit greater challenges for LVLMs, which may be attributed to the limited training resources in these specific domains. Besides, commercial models exhibit poor performance in tasks related to “people” (i.e., “race”, “age”, and “gender”). This is likely due to the additional safety training incorporated into closed-source models, which results in more conservative responses on related questions. Detailed results are provided in Appendix A.
Analysis on Intra-task. Thanks to the diverse metadata design, Dysca enables highly granular analysis within a single task. Taking animal categories as an example, we analyze the performance of LVLMs on different animals. Here, we perform same hierarchical clustering methods across 51 animal categories. As shown in the Fig. 6, LVLMs performance vary significantly across these categories. Models tend to perform poorly when applied to marine life. This could be caused by the challenges in collecting ocean-related data or possibly to the fact that human understanding of land environments is better than that of the ocean. This observation highlights the need to direct the model’s focus to oceanic domains.
4.4 The Validity of Dysca